JEN HAS A SOLUTION THAT IS 55% ISOPROPYL ALCOHOL AND A SOLUTION THAT IS 20% ISOPROPYL ALCOHOL. HOW MUCH OF EACH SHOULD SHE USE TO OBTAIN 250 ML OF A SOLUTION THAT IS 41% ISOPROPYL ALCOHOL?
Answer by Alan3354(69443)

JEN HAS A SOLUTION THAT IS 55% ISOPROPYL ALCOHOL AND A SOLUTION THAT IS 20% ISOPROPYL ALCOHOL. HOW MUCH OF EACH SHOULD SHE USE TO OBTAIN 250 ML OF A SOLUTION THAT IS 41% ISOPROPYL ALCOHOL?
--------------
f = amount of 55%
t = amount of 20%
-------
f + t = 250 (total volume)
0.55f + 0.2t = 0.41*250 (total alcohol)
You can solve them?